OpenStudy (anonymous):

solve the equation please help!! -5y-9=-(y-1)

OpenStudy (anonymous):

distribute the - to the y and the -1 on the right side of the equation making it -5y-9=-y+1 and now add 5y to the right side making it -9=4y+1 subtract the 1. -10=4y divide and simplify y=-5/2

OpenStudy (anonymous):

omg thank you :)

OpenStudy (anonymous):

You're welcome!